OpenStudy (anonymous):

Find dy/du, du/dx, and dy/dx. y = 2√u , u=9x+4

OpenStudy (anonymous):

\[\frac{dy}{du}=\frac{d}{du}[2\sqrt{u}]=\frac{1}{\sqrt{u}}\]

OpenStudy (anonymous):

\[\frac{du}{dx}=\frac{d}{dx}[9x+4]=9\]

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{dy}{du}\times \frac{du}{dx}\] so multiply them together to get the answer

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw